Home‎ > ‎Crew 160‎ > ‎

1230-XOReport

MDRS Daily Summary Report for SOL 04

Summary Title: Relay Communication & ATV Familiarization

Mission Status: We will continue to conduct our personal research as well as EVAs. We are looking to continue working with the GreenHab/plant growth studies.

Sol Activity Summary: Relay methods for constant contact with all EVA crew members and HabCom was achieved. All members are now familiarized with the ATVs. Engineer swapped batteries between two ATVs to test for functionality of equipment. Began loft water pump investigation to find any source of the problem.

Look Ahead Plan: Everyone is still excited and I will make sure that no one is overstressed. The crew has polished their communication skills as well as ATV skills, so I would like to conduct more “on foot” EVAs for now while we await our gas jug refills from the cargo shipment.

Anomalies in work:  We had a problem with the loft water pump last night (kept doing a pulse every three seconds). The anomaly stopped early in the morning. However, we found frozen/slushy water to the left of the EVA airlock outside of the HAB at the end of our EVA and an opening between the two metal sheets connected by screws outside of the HAB where the water was found.

Weather: Bright and Sunny. High visibility and low wind.

Crew Physical Status: Everyone is in great shape!

EVA: 12S-051910E-425145N

Reports to be filed:

Commander’s Report
EVA #05 Request
EVA #04 (URGENT) Summary Report
EVA #03 Summary Report
Journalist Report
HSO Report
Engineering Report

Support Requested: We have requested a refill on our gas jugs and would like an update on this matter.